OpenStudy (anonymous):

Explain the equation for this Pretty please a student increases the temperature of a 400 cm^3 balloon from 30*c to 90*c. What is the new volume?

OpenStudy (anonymous):

@campbell_st could you maybe help?

OpenStudy (anonymous):

@hartnn could you help me maybe?

OpenStudy (anonymous):

@sourwing I know youre mostly a math helper, but I cant get anyone to help over here, maybe you could?

OpenStudy (paki):

@portlandschild yeah agree with your point.... but i will help you... it's called the Charles' law... and it states that.... "if a given quantity of gas is held at a constant pressure, its volume is directly proportional to the absolute temperature"... Think of it this way... As the temperature of the gas increases, the gas molecules will begin to move around more quickly and hit the walls of their container with more force—thus the volume will increase... Keep in mind that you must use only the ""Kelvin temperature scale""" when working with temperature in all gas law formulas...! mathematically it can be written as... V1/T1 = V2/T2 =============================================== so now come to your problem.... first convert the temperatures to Kelvin... so the formula is just to add the centigrade to the kelvin absolute value... so here T1= 30C +273 = 303K... and T2 =90C +273 = 363K... now use the relation of Charles's Law... which is.... V.1 / T.1 = V.2 / T.2 or V.2 = V.1 * (T.2 / T.1) ----> equation-1 so putting the values in equation-1, we will get... V.2 = V.1 * (T.2 / T.1) V.2 = 400cm^3 * ( 363 K / 303 K ) V.2 = 400cm^3 * ( 1.19 ) V.2 = 479.20cm^3.... so this is the new volume of the balloon, if the temperature is changed from 30*c to 90*c.... hope it will be cleared now... i explained this a lot, which is against the CoC of OS, but here at OS, people mostly focus on Math, Bio, History and English section and other sections are UN-answered.... so sorry for this.... @sarah786 please have a look here too... it is useful for you too... and @portlandschild is it clear now.... :) best of luck....